# CryoBench: Diverse and challenging datasets for the heterogeneity problem in cryo-EM

### v1

@@ 5 Results
We present the results of each dataset in Figure 2-7, and we report summary statistics of volume metrics and run times in Table 1.
Method IgG-1D IgG-RL Ribosombly Tomotwin-100 Spike-MD
Mean (std) Median Time Mean (std) Med Mean (std) Med Mean (std) Med Mean (std) Med
CryoDRGN 0.366 (0.003) 0.366 0:04:00 0.349 (0.008) 0.348 0.415 (0.019) 0.415 0.321 (0.034) 0.322 0.340 (0.009) 0.340
DRGN-AI-fixed 0.366 (0.001) 0.366 0:03:28 0.355 (0.007) 0.354 0.372 (0.032) 0.374 0.206 (0.041) 0.212 0.301 (0.012) 0.304
Opus-DSD 0.34 (0.006) 0.341 0:25:51 0.346 (0.029) 0.349 0.372 (0.046) 0.382 0.256 (0.038) 0.266 0.228 (0.030) 0.242
3DFlex 0.336 (0.002) 0.336 0:28:53 0.339 (0.007) 0.339 - - - - 0.304 (0.010) 0.306
3DVA 0.351 (0.003) 0.351 0:00:29 0.341 (0.006) 0.341 0.375 (0.038) 0.372 0.095 (0.04) 0.083 0.324 (0.010) 0.325
RECOVAR 0.391 (0.001) 0.391 0:29:53 0.372 (0.008) 0.371 0.43 (0.016) 0.432 0.306 (0.093) 0.33 0.363 (0.011) 0.366
3D Class 0.297 (0.019) 0.291 0:01:15 0.309 (0.01) 0.307 0.289 (0.081) 0.288 0.046 (0.026) 0.037 0.307 (0.023) 0.308
CryoDRGN2 0.344 (0.003) 0.345 1:08:35 0.294 (0.022) 0.296 0.314 (0.09) 0.348 0.076 (0.016) 0.074 0.253 (0.037) 0.265
DRGN-AI 0.356 (0.002) 0.356 1:05:29 0.341 (0.01) 0.342 0.351 (0.052) 0.369 0.073 (0.014) 0.073 0.281 (0.009) 0.282
3D Class abinit 0.13 (0.046) 0.119 2:58:49 0.184 (0.022) 0.188 0.144 (0.036) 0.138 0.032 (0.012) 0.032 0.206 (0.010) 0.207
Table 1: AUC of Per-Conformation-FSC. Time indicates HH:MM:SS per epoch. The metrics are computed after masking out background noise. Standard deviation of all FSC-AUCs per method is given in parentheses. See Appendix C.1.3 and Figure S5 for unmasked performance. We also provide Fab-masked results for IgG-1D in Table S3 and Figure S6.
@@
IgG-1D. Figure 2c illustrates UMAP visualization of the methods’ latent spaces colored by dihedral conformations. The methods generally demonstrate compact clustering, and well preserved local neighborhoods relative to the ground truth quantified by a high neighborhood similarity and close to the equivalent region, (0,0), in the information imbalance plane (Fig. 2d). RECOVAR exhibits a latent space structure most consistent with the ground truth manifold, as shown by the embedding metrics in Figure 2d, while also achieving the highest reconstruction quality as shown in Table 1. In contrast, the rotating Fab domain is not well captured by 3DFlex (Fig. 2b), indicating limitations in effectively representing intricate conformal motions. Additionally, we provide Time per epoch for each method. For 3D Class abinit and RECOVAR, which do not have epochs, we report the whole training time. As expected, ab initio methods take significantly longer than fixed pose methods. With increasing noise levels, Figure 3a, we find that the volume metrics decrease as expected. Ab initio methods fail to reconstruct at the lowest SNR (Fig. 3b).
@@
Figure 7: Tomotwin-100 results. (a) 10 of 100 G.T. volumes and the size distribution of G.T. structures. (b) Two representative volumes for each method with the G.T. volume. (c) Latent embeddings visualized by UMAP and colored by the 100 G.T proteins. (d) Embedding comparison metrics. (e) Per-Conformation FSC plots. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation. Colors in (b), (d), and (e) correspond to methods shown in the legend. Additional results shown in Figure S19.
@@
IgG-RL. The overall performance of IgG-RL in Table 1 is lower than that of IgG-1D, attributed to the more challenging dataset design. In particular, Fab orientation inference is challenging, limiting the resolution of the moving Fab (Fig. 4c). Nonetheless, some methods, such as Opus-DSD or DRGN-AI-fixed, are able to conserve high neighborhood similarity and equivalent information with the ground truth for small radii (Fig. 4d). RECOVAR demonstrates superior performance in the FSC metric. Among ab initio methods, DRGN-AI generally outperforms the others. Despite the random orientations of the moving Fab, ab initio methods are able to align on the fixed region.
@@
Spike-MD. The high number of unique structures (46k) makes this dataset comparatively challenging with some methods modeling the continuous motion better than others by visual inspection (Fig. 5c). UMAP visualizations of the latent space reveal that methods have learned a variety of topologies, with neither of the ground truth coordinates correlating well with the different UMAP embeddings (Fig. 5d). As shown in Figure S22, the neighborhood similarity metric also shows a poor local neighborhood overlap (<35%) at small-to-intermediate radii suggesting that there is low neighborhood correlation with the ground truth, indicating that the flexibility is indeed a challenging inference task. The Per-Conformation FSC plot for Spike-MD is provided in Figure S8. Future work could explore more custom analyses and metrics for MD data, and we hope to enable future methods development connecting MD simulations with cryo-EM.
Ribosembly. Methods are largely capable of reconstructing the ground truth states (Fig. 6b,e). However, latent embeddings do not perfectly cluster each ground truth assembly state (Fig. 6c), quantified in Table 5. The neighborhood similarity ranges are intermediate, from 50-80%, for a neighborhood clustering size k corresponding to the true number of images belonging to each state (Fig. 6d). Here, the neighborhood similarity uses a rank-size embedding for ground truth heterogeneity, while information imbalance uses the voxel intensities, where similar assembly states would have a close distance. RECOVAR demonstrates better clustering, with DRGN-AI also performing well amongst ab initio methods. In contrast, both 3D Class and 3D Class abinit exhibit a mixture of ground truth states within each class, leading to lower volume accuracy as indicated in Table 1.
@@ 6 Conclusion and Future Directions
CryoBench presents: 1) five synthetic datasets containing conformational and compositional heterogeneity; 2) comprehensive analyses of ten existing state-of-the-art tools across these datasets; and 3) new quantitative metrics and qualitative visualizations to compare these baselines. Our metrics assess representation learning, reconstruction quality, and the end-to-end heterogeneous reconstruction task. Overall, we find that quantification of the reconstruction performance aligns well with qualitative observations.
@@
In future work, we aim to explore several limitations of the current benchmark. For example, we could use more complex or realistic noise statistics in image formation or simulate the joint presence of compositional and conformational heterogeneity. Through the active use of these datasets by the community, we anticipate that custom metrics designed for each type of heterogeneity will be fruitful, such as analyses of free energy landscapes and atomic-level motions in the case of MD data. Lastly, while the presented synthetic datasets provide a challenging setting for methods development, future versions of the benchmark could include non-structural sources of heterogeneity found in real cryo-EM datasets, such as junk particles and non-uniform pose distributions. We anticipate that these developments will facilitate the application of novel methods towards biological discovery on real data.
@@ B.1 Mask Generation
Certain reconstruction methods take as input a volume mask defining the region of interest. For mask generation, we first aggregated all ground truth volumes using the volume add in ChimeraX. Subsequently, we then applied the Volume Tools in CryoSPARC. Specifically, for IgG-1D, IgG-RL, and Ribosembly, the Dilation radius (pix) and Soft padding width (pix) were set at 8 and 5, respectively. For Tomotwin-100, these parameters were adjusted to a Dilation radius (pix) of 5 and a Soft padding width (pix) of 3. For Spike-MD, we take the union of all binarized volumes and use the cryoDRGN gen_mask command with a dilation of 25 Å and soft cosine edge of 15 Å. Masks for each dataset are shown in Figure S3b.
@@ B.2 CryoDRGN, CryoDRGN2
B.3 DRGN-AI, DRGN-AI-fixed
DRGN-AI [47] is a deep generative network-based method, inspired by cryoDRGN. DRGN-AI uses both HPS and stochastic gradient descent in pose estimation, while utilizing a differential lookup table instead of an encoder network to encode the pose and conformational latent variable information. We denote the fixed pose mode of operation with “DRGN-AI-fixed” and ab initio with “DRGN-AI.”
We trained DRGN-AI and DRGN-AI-fixed using the official PyTorch implementation [^2] https://github.com/ml-struct-bio/drgnai, version 0.2.2b0. We used the default settings with the z-dimension set to 4 and the total number of training epochs set to 100. We used one A100 GPU for training.
@@ B.10 Information Imbalance
B.11 Training Time
Experiments for the training times reported in Table 1 were conducted on a single A100 GPU with 512G of CPU memory.
@@ C.1.2 Sensitivity to Structural Heterogeneity
We first validate that our volume comparison metric, the area under the Fourier Shell Correlative curve (AUC-FSC), is sensitive to structural differences between ground truth volumes. In Figure S4, we compute AUC-FSC between different ground truth volumes for the IgG-1D dataset. The FSC-AUC between a given volume and all other volumes reaches its highest point (0.5) at the ground truth index and is unimodal across the ground truth conformational coordinate (note that IgG-1D describes a circular 1D motion and thus the x-axis is cyclic) (Fig. S4a). Figure S4b summarizes the all-to-all comparison.
@@
Method CryoDRGN DRGN-AI-fixed Opus-DSD 3DFlex 3DVA RECOVAR 3D Class
@@
Method CryoDRGN2 DRGN-AI 3D Class abinit
@@ C.2 Volume FSC metrics for heterogeneous reconstruction
To compute Per-Conformation FSC, we sample a representative volume $V^{\prime}$ for each ground truth conformation $V$ by averaging the latent embeddings for each particle image of $V$. We present FSC curves averaged across all ground truth conformations for all methods for IgG-1D, IgG-RL, Ribosembly, and Tomotwin-100 in Figures 2, 4, 6, 7, respectively. Per-Conformation FSC curves for the Spike-MD dataset are given in Figure S8. We additionally show all 100 FSC curves for the IgG-1D dataset for each method in Figure S7.
@@
Mean (std) Median Mean (std) Med Mean (std) Med Mean (std) Med
@@
Sample FSC. While Per-Conformation FSC is the main metric we report, it may leverage information from the ground truth clustering when computing representative latents/volumes. As an alternative, we also generate volumes after unsupervised clustering of the latent embeddings to yield representative samples (without any use of G.T. information). However, as there is then no obvious correspondence between the sampled volume and its G.T. conformation, we compute the AUC-FSC for each sampled volume against all G.T. volumes and take the max value. We call this quantification the Sample FSC and Table S5 shows the AUC-FSC values.
@@
Figure S10: Per-image FSC. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation.
Per-Image FSC. Finally, we use Per-image FSC as a metric that jointly assesses conformation estimation and reconstruction quality. Here, we use a set of $N$ images from the dataset (either sampled at random or uniformly for each conformation). For each image, we reconstruct an associated volume at its inferred latent coordinate. We then compute the FSC AUC for the reconstructed volume to the image’s ground truth volume. Thus, unlike Per-Conformation FSC and Sample FSC, Per-Image FSC assesses how well each method has reconstructed the distribution of volumes. We use 100 images for each dataset to compute Per-Image FSC with one random image per ground truth conformation. Figure S10 provides Per-image FSC curves for each method across all datasets and Table S5 shows the AUC-FSC values.
Method IgG-1D IgG-RL Ribosombly Tomotwin-100
Mean (std) Median Mean (std) Med Mean (std) Med Mean (std) Med
CryoDRGN 0.351 (0.028) 0.356 0.331 (0.016) 0.333 0.412 (0.023) 0.415 0.316 (0.046) 0.321
DRGN-AI-fixed 0.364 (0.002) 0.364 0.348 (0.012) 0.350 0.372 (0.032) 0.375 0.202 (0.044) 0.207
Opus-DSD 0.335 (0.026) 0.339 0.343 (0.016) 0.346 0.362 (0.083) 0.382 0.237 (0.049) 0.251
3DFlex 0.335 (0.003) 0.335 0.337 (0.007) 0.337 - - - -
3DVA 0.349 (0.004) 0.350 0.333 (0.014) 0.335 0.375 (0.038) 0.375 0.088 (0.04) 0.077
RECOVAR 0.386 (0.005) 0.388 0.363 (0.011) 0.363 0.429 (0.018) 0.432 0.258 (0.109) 0.254
3D Class 0.297 (0.019) 0.291 0.309 (0.01) 0.307 0.289 (0.081) 0.288 0.046 (0.026) 0.037
CryoDRGN2 0.32 (0.062) 0.342 0.301 (0.03) 0.306 0.341 (0.059) 0.356 0.076 (0.016) 0.072
DRGN-AI 0.351 (0.01) 0.352 0.329 (0.028) 0.333 0.341 (0.083) 0.367 0.072 (0.015) 0.072
3D Class abinit 0.13 (0.046) 0.119 0.184 (0.022) 0.188 0.144 (0.036) 0.138 0.032 (0.012) 0.031
Table S6: AUC of Per-image FSC. The metrics are computed after masking out background noise. Standard deviation of all FSC-AUCs per method is given in parentheses.

### v2

@@ 5 Results
We present the benchmarking results for each dataset in Figures 2-7, and we report summary statistics of volume metrics in Table 1.
Method IgG-1D IgG-RL Ribosembly Tomotwin-100 Spike-MD
Mean (std) Median Mean (std) Med Mean (std) Med Mean (std) Med Mean (std) Med
CryoDRGN 0.351 (0.028) 0.356 0.331 (0.016) 0.333 0.412 (0.023) 0.415 0.316 (0.046) 0.321 0.340 (0.009) 0.340
CryoDRGN-AI-fixed 0.364 (0.002) 0.364 0.348 (0.012) 0.350 0.372 (0.032) 0.375 0.202 (0.044) 0.207 0.301 (0.012) 0.303
Opus-DSD 0.335 (0.026) 0.339 0.343 (0.016) 0.346 0.362 (0.083) 0.382 0.237 (0.049) 0.251 0.229 (0.027) 0.242
3DFlex 0.335 (0.003) 0.335 0.337 (0.007) 0.337 - - - - 0.304 (0.011) 0.306
3DVA 0.349 (0.004) 0.350 0.333 (0.014) 0.335 0.375 (0.038) 0.375 0.088 (0.04) 0.077 0.324 (0.010) 0.323
RECOVAR 0.386 (0.005) 0.388 0.363 (0.011) 0.363 0.429 (0.018) 0.432 0.258 (0.109) 0.254 0.362 (0.011) 0.365
3D Class 0.297 (0.019) 0.291 0.309 (0.01) 0.307 0.289 (0.081) 0.288 0.046 (0.026) 0.037 0.307 (0.023) 0.308
CryoDRGN2 0.32 (0.062) 0.342 0.301 (0.03) 0.306 0.341 (0.059) 0.356 0.076 (0.016) 0.072 0.245 (0.042) 0.260
CryoDRGN-AI 0.351 (0.01) 0.352 0.329 (0.028) 0.333 0.341 (0.083) 0.367 0.072 (0.015) 0.072 0.279 (0.017) 0.281
3D Class abinit 0.13 (0.046) 0.119 0.184 (0.022) 0.188 0.144 (0.036) 0.138 0.032 (0.012) 0.031 0.206 (0.009) 0.208
Table 1: Area under the Per-Image FSC Curve. FSC curves are computed after masking out background noise. See Appendix C.1.3 and Figure S6 for unmasked performance. We also provide results for IgG-1D with a mask only around the FAb in Table S3, Table S4, and Figure S7.
@@
IgG-1D. All methods generally demonstrate interpretable latent spaces (Fig. 2c) and well-preserved local neighborhoods relative to the ground truth quantified by a high neighborhood similarity and close to the equivalent region, (0,0), in the information imbalance plane (Fig. 2d). RECOVAR exhibits a latent space structure most consistent with the ground truth manifold, as shown by the embedding metrics in Figure 2d, while also achieving the highest reconstruction quality as shown in Table 1. In contrast, the rotating Fab domain is not well captured by 3DFlex (Fig. 2b), indicating limitations in effectively representing intricate conformal motions. With increasing noise levels, we find that the volume metrics decrease as expected (Fig. 3a). Ab initio methods fail to reconstruct at the lowest SNR (Fig. 3b).
@@
Figure 7: Tomotwin-100 results. a) Dataset design. Density maps for 10 out of 100 ground truth volumes and a histogram of molecular weights. b) Representative reconstructed and ground truth (G.T.) volumes. c) Latent embeddings visualized by UMAP and colored by the 100 G.T. molecular weight. d) Latent embedding analysis by neighborhood similarity and information imbalance. e) Per-Image FSC curves. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation. Colors in (b), (d), and (e) correspond to methods shown in the legend. Additional results shown in Figure S21.
@@
IgG-RL. The overall performance of IgG-RL in Table 1 is lower than that of IgG-1D, attributed to the more challenging dataset design. In particular, Fab orientation inference is challenging, limiting the resolution of the moving Fab (Fig. 4c). Nonetheless, some methods, such as Opus-DSD or CryoDRGN-AI-fixed, are able to conserve high neighborhood similarity and equivalent information with the ground truth for small radii (Fig. 4d). RECOVAR demonstrates superior performance in the FSC metric. Among ab initio methods, CryoDRGN-AI generally outperforms the others. Despite the random orientations of the moving Fab, ab initio methods are able to align on the fixed region.
@@
Spike-MD. The high number of unique structures (46k) makes this dataset comparatively challenging with some methods modeling the continuous motion better than others by visual inspection (Fig. 5c). UMAP visualizations of the latent space reveal that methods have learned a variety of topologies, with neither of the ground truth coordinates correlating well with the different UMAP embeddings (Fig. 5d). As shown in Figure S24, the neighborhood similarity metric also shows a poor local neighborhood overlap (<35%) at small-to-intermediate radii suggesting that there is low neighborhood correlation with the ground truth, indicating that the flexibility is indeed a challenging inference task. The Per-Image FSC plot for Spike-MD is provided in Figure S12. Future work could explore more custom analyses and metrics for MD data, and we hope to enable future methods development connecting MD simulations with cryo-EM.
@@
Ribosembly. Methods are largely capable of reconstructing the ground truth states (Fig. 6b,e). However, latent embeddings do not perfectly cluster each ground truth assembly state (Fig. 6c), quantified in Table 5. The neighborhood similarity ranges are intermediate, from 50-80%, for a neighborhood clustering size k corresponding to the true number of images belonging to each state (Fig. 6d). Here, the neighborhood similarity uses a rank-size embedding for ground truth heterogeneity, while information imbalance uses the voxel intensities, where similar assembly states would have a close distance. RECOVAR demonstrates better clustering, with CryoDRGN-AI also performing well amongst ab initio methods. In contrast, both 3D Class and 3D Class abinit exhibit a mixture of ground truth states within each class, leading to lower volume accuracy as indicated in Table 1.
@@ 6 Conclusion and Future Directions
CryoBench presents: 1) five synthetic datasets containing conformational and compositional heterogeneity; 2) comprehensive analyses of ten existing state-of-the-art tools across these datasets; and 3) quantitative metrics and qualitative visualizations to compare these baselines. Our metrics assess representation learning, reconstruction quality, and the end-to-end heterogeneous reconstruction task. Overall, we find that quantification of the reconstruction performance aligns well with qualitative observations.
@@
In future work, we aim to explore several limitations of the current benchmark. For example, we could use more complex or realistic noise statistics in image formation or simulate the joint presence of compositional and conformational heterogeneity. To explore realistic noise statistics in image formation, we tested cisTEM’s (Computational Imaging System for Transmission Electron Microscopy) [51] multislice simulator, which considers artifacts from sample thickness, noise from explicitly modeled water molecules, and the effects of sample motion and radiation damage as a function of electron dose. However, we note that this simulator requires several minutes to generate each image. Since the images it produced and the performance of each method were qualitatively similar to those generated using our simpler Gaussian noise model, we did not follow this procedure for the benchmark datasets.
Through the active use of these datasets by the community, we anticipate that custom metrics designed for each type of heterogeneity will be fruitful, such as analyses of free energy landscapes and atomic-level motions in the case of MD data. Lastly, while the presented synthetic datasets provide a challenging setting for methods development, future versions of the benchmark could include non-structural sources of heterogeneity found in real cryo-EM datasets, such as junk particles and non-uniform pose distributions. We anticipate that these developments will facilitate the application of novel methods towards biological discovery on real data.
@@ A.2 Generating IgG-RL
Figure S1: Additional details for IgG-RL dataset design. a) The dihedral angles of a linking peptide are chosen to randomly orient one arm of the IgG molecule in 100 unique conformations. b) The distribution of the 100 structures across coordinates can be characterized by the angle and distance between the rotating arm and the rigid arms.
@@ B.1 Mask Generation
Some reconstruction methods require as input a volume mask defining the region of interest. For mask generation, we first aggregated all ground truth volumes using the volume add in ChimeraX. Subsequently, we then applied the Volume Tools in CryoSPARC. Specifically, for IgG-1D, IgG-RL, and Ribosembly, the Dilation radius (pix) and Soft padding width (pix) were set at 8 and 5, respectively. For Tomotwin-100, these parameters were adjusted to a Dilation radius (pix) of 5 and a Soft padding width (pix) of 3. For Spike-MD, we take the union of all binarized volumes and use the cryoDRGN gen_mask command with a dilation of 25 Å and soft cosine edge of 15 Å. Masks for each dataset are shown in Figure S4b.
@@ B.2 CryoDRGN, CryoDRGN2
B.3 CryoDRGN-AI, CryoDRGN-AI-fixed
CryoDRGN-AI [47] is a deep generative network-based method, inspired by cryoDRGN. CryoDRGN-AI uses both HPS and stochastic gradient descent in pose estimation, while utilizing a differential lookup table instead of an encoder network to encode the pose and conformational latent variable information. We denote the fixed pose mode of operation with “CryoDRGN-AI-fixed” and ab initio with “CryoDRGN-AI.”
We trained CryoDRGN-AI and CryoDRGN-AI-fixed using the official PyTorch implementation [^2] https://github.com/ml-struct-bio/drgnai, version 0.2.2b0. We used the default settings with the z-dimension set to 4 and the total number of training epochs set to 100. We used one A100 GPU for training.
@@ C.1.2 Sensitivity to Structural Heterogeneity
We first validate that our volume comparison metric, the area under the Fourier Shell Correlative curve ($\text{FSC}_{\text{AUC}}$), is sensitive to structural differences between ground truth volumes. In Figure S5, we compute $\text{FSC}_{\text{AUC}}$ between different ground truth volumes for the IgG-1D dataset. The FSC-AUC between a given volume and all other volumes reaches its highest point (0.5) at the ground truth index and is unimodal across the ground truth conformational coordinate (note that IgG-1D describes a circular 1D motion and thus the x-axis is cyclic) (Fig. S5a). Figure S5b summarizes the all-to-all comparison.
@@
Method CryoDRGN CryoDRGN-AI-fixed Opus-DSD 3DFlex 3DVA RECOVAR 3D Class
@@
Method CryoDRGN2 CryoDRGN-AI 3D Class abinit
@@ C.2 Volume FSC metrics for heterogeneous reconstruction
Figure S8: Per-Conformation FSC. Each curve shows the average FSC curve across all conformations with error bars indicating the standard deviation.
@@
For each ground truth conformation $V$, we compute a corresponding latent coordinate by averaging all latent embeddings $\bar{z}$ of images generated by $V$. To remain on the data manifold, we identify the closest latent coordinate $z^{*}$ to $\bar{z}$. Using this coordinate, we reconstruct the associated volume and evaluate its $\text{FSC}_{\text{AUC}}$ against the ground truth volume. Figure S8 and S9 provides Per-Conformation FSC curves for each method across all datasets and Table S5 shows the $\text{FSC}_{\text{AUC}}$ values. We additionally show all 100 FSC curves for the IgG-1D dataset for each method in Figure S10.
@@
Mean (std) Med Mean (std) Med Mean (std) Med Mean (std) Med
@@
Sample FSC. While Per-Image FSC is the main metric we report, it may leverage information from the ground truth clustering when computing representative latents/volumes. As an alternative, we also generate volumes after unsupervised clustering of the latent embeddings to yield representative samples (without any use of G.T. information). However, as there is then no obvious correspondence between the sampled volume and its G.T. conformation, we compute the $\text{FSC}_{\text{AUC}}$ for each sampled volume against all G.T. volumes and take the max value. We call this quantification the Sample FSC and Table S5 shows the $\text{FSC}_{\text{AUC}}$ values.
@@
Per-Image FSC. Finally, we use Per-Image FSC as a metric that jointly assesses conformation estimation and reconstruction quality. Here, we use a set of $N$ images from the dataset (either sampled at random or uniformly for each conformation). For each image, we reconstruct an associated volume at its inferred latent coordinate. We then compute the FSC AUC for the reconstructed volume to the image’s ground truth volume. Thus, unlike Per-Conformation FSC and Sample FSC, Per-Image FSC assesses how well each method has reconstructed the distribution of volumes. We use 100 images for each dataset to compute Per-Image FSC with one random image per ground truth conformation.
Figure S12: Per-Image FSC for Spike-MD. Each curve shows the average FSC curve across 100 sampled images with error bars indicating the standard deviation.
Method IgG-1D IgG-RL Ribosembly Tomotwin-100 Spike-MD
Mean (std) Med Mean (std) Med Mean (std) Med Mean (std) Med Mean (std) Med
CryoDRGN 0.366 (0.003) 0.366 0.349 (0.008) 0.348 0.415 (0.019) 0.415 0.321 (0.034) 0.322 0.340 (0.009) 0.340
CryoDRGN-AI-fixed 0.366 (0.001) 0.366 0.355 (0.007) 0.354 0.372 (0.032) 0.374 0.206 (0.041) 0.212 0.301 (0.012) 0.304
Opus-DSD 0.34 (0.006) 0.341 0.346 (0.029) 0.349 0.372 (0.046) 0.382 0.256 (0.038) 0.266 0.228 (0.030) 0.242
3DFlex 0.336 (0.002) 0.336 0.339 (0.007) 0.339 - - - - 0.304 (0.010) 0.306
3DVA 0.351 (0.003) 0.351 0.341 (0.006) 0.341 0.375 (0.038) 0.372 0.095 (0.04) 0.083 0.324 (0.010) 0.325
RECOVAR 0.391 (0.001) 0.391 0.372 (0.008) 0.371 0.43 (0.016) 0.432 0.306 (0.093) 0.33 0.363 (0.011) 0.366
3D Class 0.297 (0.019) 0.291 0.309 (0.01) 0.307 0.289 (0.081) 0.288 0.046 (0.026) 0.037 0.307 (0.023) 0.308
CryoDRGN2 0.344 (0.003) 0.345 0.294 (0.022) 0.296 0.314 (0.09) 0.348 0.076 (0.016) 0.074 0.253 (0.037) 0.265
CryoDRGN-AI 0.356 (0.002) 0.356 0.341 (0.01) 0.342 0.351 (0.052) 0.369 0.073 (0.014) 0.073 0.281 (0.009) 0.282
3D Class abinit 0.13 (0.046) 0.119 0.184 (0.022) 0.188 0.144 (0.036) 0.138 0.032 (0.012) 0.032 0.206 (0.010) 0.207
Table S6: AUC of Per-Conformation-FSC. The metrics are computed after masking out background noise. Standard deviation of all FSC-AUCs per method is given in parentheses.
We present FSC curves averaged across all ground truth conformations for all methods for IgG-1D, IgG-RL, Ribosembly, and Tomotwin-100 in Figures 2, 4, 6, 7, respectively. FSC curves for Spike-MD are shown in Figure S12.